OpenStudy (anonymous):

A kangaroo jumps off the ground with an initial velocity of 18 feet per second. a) Write an equation that gives the height (in feet) of the kangaroo as a function of time (in seconds) since it jumps. b) After how many seconds does that kangaroo land on the ground?

OpenStudy (anonymous):

@freckles can you help me with this?

OpenStudy (freckles):

so we know we have a parabola type graph

OpenStudy (freckles):

we are given the initial height is 0 ft and the initial velocity is 18 ft/sec

OpenStudy (anonymous):

ok

OpenStudy (freckles):

h(t)=at^2+bt+c We have that c is the initial height b is the initial velocity and I forgot with a is a is usually a set number a is -16 I think

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so would the formula be: y=-16t^2+18t?

OpenStudy (freckles):

that a is the number that has to do with gravity yeah is it -16 is you are dealing with feet and -4.9 if you are dealing with meters

OpenStudy (freckles):

and we definitely have feet

OpenStudy (freckles):

yep

OpenStudy (anonymous):

ok

OpenStudy (freckles):

the next question on the ground means the height is 0 correct?

OpenStudy (freckles):

so find t for when y is 0

OpenStudy (freckles):

that is solve 0=-16t^2+18t for t

OpenStudy (anonymous):

is it 9/8?

OpenStudy (freckles):

yep you have the the solution t=0 or t=9/8... t=0 sec is right before the kangaroo lifts off and t=1.125 or 9/8 sec is when you have the kangaroo lands back on the ground after his hop

OpenStudy (anonymous):

ok awesome... thanks again you r really smart

OpenStudy (freckles):

and of course yes you are looking for the 1.125 solution since it asks after how many sec does he land back on the ground

OpenStudy (freckles):

more examples like this problem can be found here: http://www.purplemath.com/modules/quadprob.htm

OpenStudy (freckles):

anyways peace and have fun

OpenStudy (anonymous):

OK GREAT! THX